An object is released from a height.How to find the distance travelled by an object if distance travelled in first 3 seconds is equal to the distance travelled in the last second are equal. explain

Grade:11

1 Answers

vishal akula
6 years ago
let us assume the situation as ball released from some height
a=g=10m/s2 u=0 let the time taken to reach ground be n
==> in first 3 seconds
 dis(s)=1/2at^2= 45m---{1}
dis convered in nth second formula=s=1/2a(2n-1)-- {2}      
5(2n-1)=45
n=5 total distance=5(5)(5)=125